# Automorphism group schemes of lattice vertex operator algebras

### v1

@@ Abstract.
Given a positive definite even lattice and a commutative ring, there is a standard construction of a lattice vertex algebra over the commutative ring, and this is promoted to a vertex operator algebra when the determinant of the lattice is invertible. We describe the groups of automorphisms of these vertex algebras and vertex operator algebras as affine group schemes, showing in particular that each is an extension of an explicitly described split reductive group of ADE type by the outer automorphism group of the lattice.
@@ Introduction
The statement of Dong-Nagatomo’s theorem is that if $L$ is a positive-definite even lattice, then the automorphism group of the lattice vertex operator algebra $V_{L,\mathbb{C}}$ is a product of a certain Lie group $N$ with a finite group $O(\hat{L})$. They also show that $N\cap O(\hat{L})$ contains the 2-torsion group $\operatorname{Hom}(L,\pm 1)$. Our result (Theorem 4.4.1) replaces $N$ with a split reductive algebraic group $G_{L}$ defined over $\mathbb{Z}$, and the finite group $O(\hat{L})$ with a finite flat group scheme $O(\tilde{L})$. Naturally, the complex analytification of $G_{L}$ is isomorphic to $N$, and $O(\tilde{L})(\mathbb{C})\cong O(\hat{L})$. We also determine the intersection $G_{L}\cap O(\tilde{L})$ precisely, identifying it with the “Tits group” of $G_{L}$, and use this to show that the quotient $\operatorname{Aut}V_{L}/G_{L}$ is isomorphic to the outer automorphism group $\operatorname{Aut}L/W_{L}$ of the lattice, i.e., the quotient of all orthogonal transformations by the subgroup generated by reflections.
In addition to these generalizations, we remove the hypothesis that $V_{L}$ has a conformal vector, instead considering homogeneous vertex algebra automorphisms. This effectively means we allow ourselves to consider commutative rings $R$ where the determinant of $L$ is not invertible. The fact that we still get a uniform description of the automorphism group is somewhat surprising, since $V_{L}$ is no longer a simple vertex algebra when $\det L=0$ in $R$. In particular, the full ring of endomorphisms can become much larger than normally expected over such $R$. One concrete example is the $A_{2}$ lattice over $\mathbb{F}_{3}$, whose vertex algebra has a simple quotient with automorphism group $G_{2}(3)$: see [Griess-Lam 2013], Proposition 6.1 and Theorem 6.3  [^1] We thank C. H. Lam for bringing this example to our attention..
Viewed from a suitable distance, our proof is more or less a translation of Dong-Nagatomo’s proof, with some necessary adjustments. One such adjustment lies in the construction of $G_{L}$. Dong and Nagatomo construct $N$ by simply exponentiating the weight 1 Lie algebra of $V_{L,\mathbb{C}}$, but reductive algebraic groups in general cannot be constructed by exponentiating a Lie algebra. Instead, we construct $G_{L}$ by describing its torus $T=\operatorname{\underline{Hom}}(L,\mathbb{G}_{m})$ separately from its one-dimensional root subgroups $G_{\alpha}$, then showing that these subgroups satisfy the appropriate relations.
The main sticking point for us was the conjugacy of tori: over an algebraically closed field, all Cartan subgroups of a smooth affine group scheme are conjugate to each other, but this doesn’t hold for general commutative rings. Since the maximal tori we consider are split, there are fairly strong replacement results: Proposition 6.2.11 in [Conrad 2014] asserts that if $G$ is split reductive over a nonempty connected scheme $S$ with trivial Picard group, then all split maximal tori are conjugate. It seems possible to turn this result into a solution, but we found it more straightforward to use étale-local conjugacy of all maximal tori. Using the étale sheaf property of affine schemes, we find that this is sufficient to show that all automorphisms are covered.
@@
In section 2, we introduce some basic theory of reductive group schemes, and define the group $G_{L}$. It is a split reductive group over $\mathbb{Z}$, with connected fibers, and it acts faithfully on $V_{L}$.
In section 3, we define the finite flat group scheme $O(\tilde{L})$, and show that it acts faithfully on $V_{L}$.
@@ Acknowledgements
This work was supported by JSPS Kakenhi Grant-in-Aid for Young Scientists (B) 17K14152, and JSPS Kakenhi Grant-in-Aid for Scientific Research (C) 22K03264
@@ 1.1. Vertex algebras
    $[u_{m},v_{n}]=\sum_{k\geq 0}\binom{m}{k}(u_{k}v)_{m+n-k},$
    or its equivalent $[u_{m},Y(v,z)]=\sum_{k\geq 0}\binom{m}{k}Y(u_{k}v,z)z^{m-k}$
@@
  • (translation-covariance) $[T^{(m)},Y(u,z)]=\sum_{i=1}^{m}\partial^{(i)}Y(u,z)T^{(m-i)}$, where $T^{(m)}:V\to V$ is the linear map $u\mapsto u_{-n-1}\mathbf{1}$.
@@
    $L_{0}$ acts semisimply, i.e., each vector is a finite sum of eigenvectors.
@@
    Each $V_{n}$ is a finite projective $R$-module.
@@
    For all $m,n\in\mathbb{Z}$, $[L_{m},L_{n}]=(m-n)L_{m+n}+\binom{m+1}{3}c\operatorname{id}_{V}$.
@@
The assumption that graded pieces are finite and projective implies the image of any finite generating set lies in some finite projective module, so we obtain an embedding in a product of general linear groups. The condition that vertex algebra multiplication is respected is a Zariski closed condition, so $\operatorname{\underline{Aut}}_{V/R}$ is represented by a closed affine group subscheme of some $GL_{n,R}$. ∎
@@ 1.2. Lattice vertex algebras
Let $L$ be an integer lattice, i.e., a finitely generated free abelian group equipped with a nondegenerate integer-valued symmetric bilinear form. We define the rational Heisenberg Lie algebra $H_{L}$ on $L$ to be the $\mathbb{Q}$-vector space $(L\otimes\mathbb{Q})[t,t^{-1}]\oplus\mathbb{Q}K$ with bracket given by $[t^{i},t^{j}]=i\delta_{i+j,0}$ and $[t_{i},K]=0$. This is a central extension of the abelian Lie algebra $(L\otimes\mathbb{Q})[t,t^{-1}]$ by the one dimensional Lie algebra $\mathbb{Q}K$. We write $H_{L}^{\geq 0}$ for the abelian subalgebra $(L\otimes\mathbb{Q})[t]\oplus\mathbb{C}K$, and for any $a\in L\otimes\mathbb{Q}$ and $r\in\mathbb{Q}$, we write $\mathbb{Q}_{a,r}$ for the 1-dimensional representation on which the subalgebra $t(L\otimes\mathbb{Q})[t]$ acts trivially, $(L\otimes\mathbb{Q})$ acts by inner product with $a$, and $K$ acts by the scalar $r$. Then, we write $\pi^{L}_{a,\mathbb{Q}}$ for the induced module $\operatorname{Ind}_{H_{L}^{\geq 0}}^{H_{L}}\mathbb{Q}_{a,1}$, and identify this space with $\operatorname{Sym}_{\mathbb{Q}}(t^{-1}(L\otimes\mathbb{Q})[t^{-1}])\otimes\mathbb{Q}e_{a}$ for a distinguished vector $e_{a}$. For any $b\in L\otimes\mathbb{Q}$ and $n\in\mathbb{Z}$, we define $b(n)$ to be the operator
@@
([Frenkel-Lepowsky-Meurman 1988] chapter 5) Let $L$ be an even lattice, i.e., an integer lattice where $(a,a)\in 2\mathbb{Z}$ for all $a\in L$. We define the double cover $\hat{L}$ to be the unique (up to isomorphism) central extension of $L$ by ±$\pm 1$ such that the commutator $[a,b]$ is given by $(-1)^{(\bar{a},\bar{b})}$, where $\bar{a},\bar{b}\in L$ are the images of $a,b\in\hat{L}$ under the covering homomorphism $\hat{L}\to L$. We will write $e:L\to\hat{L}$, given by $\alpha\mapsto e_{\alpha}$, to denote a set-theoretic section of the covering homomorphism. The twisted group ring $\mathbb{Z}\{L\}$ of $L$ is the quotient of $\mathbb{Z}[\hat{L}]$ by the ideal generated by $\kappa+1$, where $\kappa$ is the nontrivial element in $\hat{L}$ that maps to $0\in L$. We write $\iota:\hat{L}\to\mathbb{Z}\{L\}$, given by $a\mapsto a\otimes 1$ for the multiplicative map. We note that there is an isomorphism $\mathbb{Z}[L]\to\mathbb{Z}\{L\}$ of abelian groups taking the basis vector $e^{\alpha}$ to $\iota(e_{\alpha})$, but this is in general not a ring homomorphism.
@@ 2.1. Fundamentals
We review the parts of the theory of reductive groups that are necessary for us. A basic reference is section 1 of [Conrad 2014].
@@
By [SGA3] Exp. XIV Corollary 3.20, maximal tori of smooth group $R$-schemes exist Zariski locally on $\operatorname{Spec}R$. By [SGA3] Exp. XXII Corollary 5.3.11, any two maximal tori (of an fpqc locally split fiberwise connected reductive group) are conjugate étale locally. This fact will play an essential role in our analysis of $\operatorname{Aut}V_{L}$.
@@
An affine group scheme over a field is reductive if it is reduced, of finite type, and its maximal connected unipotent normal subgroup is trivial (here, a group is unipotent if it admits an increasing filtration whose quotients are subgroups of the additive group $\mathbb{G}_{a}$, the group scheme which takes a ring to its underlying additive group). Given a commutative ring $R$, a reductive group scheme over $R$ is a group scheme $G\to\operatorname{Spec}R$ whose structure morphism is finite type, flat, and affine, and whose geometric fibers are reductive in the previous sense.
@@
A root datum is reduced if for any root $a\in\Phi$, the only scalar multiples that are roots are $a,-a$. A root datum is semisimple if the saturation of $\mathbb{Z}\Phi$ in $X$ is all of $X$, or equivalently if $\mathbb{Q}\Phi=X\otimes\mathbb{Q}$.
@@
By [SGA3] Exp. XIV Corollary 3.20, for any smooth reductive group scheme $G$ over a commutative ring $R$ with connected fibers, and any point $x\in\operatorname{Spec}R$, there is a Zariski open neighborhood $U$ of $x$ such that $G_{U}$ has a maximal $U$-torus. Since such a torus is étale-locally split, we consider triples $(G,T,M)$, where $T=D_{R}(M)$ is split. Then, the adjoint $T$-action on the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ endows the vector bundle $\mathfrak{g}$ with an $M$-grading, splitting into a direct sum of $\operatorname{Lie}(T)$ and line bundles $\mathfrak{g}_{a}$. The corresponding sections $a\in M$ are called roots of the Lie algebra.
@@
$G$ is smooth and affine over $\mathbb{Z}$ with connected fibers, and its Lie algebra has distinguished Cartan subalgebra $X$ and roots given by $\Phi$.
@@
(basic properties of our group) $G_{L}$ is a smooth affine group scheme over $\mathbb{Z}$ with connected fibers. The Lie algebra of $G_{L}$ is the split reductive Lie algebra over $\mathbb{Z}$ given by the root datum $(X,\Phi,X^{\vee},\Phi^{\vee})$, and in particular is isomorphic to the weight 1 subspace $(V_{L})_{1}$, where the Cartan subalgebra $X$ is identified with the weight 1 subspace of $\pi^{L}_{0}$, with roots given by $\Phi$. The semisimple group $D(G_{L})$ is of ADE type (i.e., it is simply laced), and there is a central isogeny $Z\times D(G_{L})\to G_{L}$ with finite central kernel.
@@ 3. Symmetries of the Mewtwo cover
The symmetry group of this cover is then an extension $O(\tilde{L})$ of $O(L)$ by $\operatorname{Hom}(L,\mu_{2})\cong\mu_{2}^{\operatorname{rank}L}$. We show that $O(\tilde{L})$ acts faithfully on $V_{L}$.
@@ 3.2. Orthogonal transformations
Let $(\tilde{A},\pi)$ be the $\mu_{2}$-cover of a free abelian group $A$ with quadratic form $q:A\to\mathbb{Z}$. An orthogonal transformation of $(\tilde{A},\pi)$ over $S$ is a group scheme automorphism $\phi$ of $\tilde{A}_{S}$ such that there exists an automorphism $\bar{\phi}$ of $A$ that preserves $q$, satisfying $\pi\circ\phi=\bar{\phi}\circ\pi$. The functor that takes an $R$-algebra $S$ to the group of orthogonal transformations of $(\tilde{A},\pi)$ over $S$ is written $O(\tilde{A})$, or $O(\tilde{A},\pi)$.
@@
We now consider the kernel. We first note that for any $\lambda:\underline{A}_{S}\to\mu_{2,S}$, the map $\lambda^{*}:\tilde{A}(S)\to\tilde{A}(S)$ is nontrivial if $\lambda$ is nontrivial, and induces the trivial map $\operatorname{Spec}S\to O(A,q)$, so $\operatorname{\underline{Hom}}(\underline{A},\mu_{2})$ lies in the kernel of the homomorphism. If $g\in O(\tilde{A},\pi)(S)$ induces the trivial map $\operatorname{Spec}S\to O(A,q)$, then $\pi\circ g=\pi$ as $S$-group scheme homomorphisms $\tilde{A}_{S}\to\underline{A}_{S}$. For any $x\in A$, we have a subscheme $\hat{x}\in\underline{A}_{S}$ corresponding to $x$ that is isomorphic to $\operatorname{Spec}S$, and $\pi^{-1}(\hat{x})$ is a $\mu_{2,S}$-torsor. The restriction of $g$ to $\pi^{-1}(\hat{x})$ is then multiplication by some element of $\mu_{2,S}(S)$. The group law on $\tilde{A}_{S}$ implies $g$ is uniquely determined by the restriction of $g$ to $\pi^{-1}(\{\hat{e}_{1},\ldots,\hat{e}_{r}\})$, where $e_{1},\ldots,e_{r}$ is a basis of $A$, and any such choice of $r$ elements of $\mu_{2,S}(S)$ corresponds to a unique element of $\operatorname{\underline{Hom}}(\underline{A},\mu_{2})(S)$.
@@
If $L$ is positive-definite, then $O(\tilde{L})$ is an affine group scheme of finite type.
@@
Given a basis of $L$, let $d$ be the maximal norm of vectors in this basis, and let $n$ be the number of vectors in $L$ of length at most $d$. Then $O(\tilde{L})$ is a subgroup of $GL_{n}$, since it acts faithfully on the linear space whose basis is the set of these vectors. ∎

### v2

# Automorphism group schemes of lattice vertex algebras

@@ Abstract.
Given a positive definite even lattice and a commutative ring, there is a standard construction of a lattice vertex algebra over the commutative ring, and it admits a natural grading by non-negative integers. We describe the groups of automorphisms of these graded vertex algebras as affine group schemes, showing in particular that each is an extension of an explicitly described split reductive group of ADE type by the outer automorphism group of the lattice.
@@ Introduction
[…]
In addition to these generalizations, we remove the hypothesis that $V_{L}$ has a conformal vector, instead considering homogeneous vertex algebra automorphisms. This effectively means we allow ourselves to consider commutative rings $R$ where the determinant of $L$ is not invertible. The fact that we still get a uniform description of the automorphism group is somewhat surprising, since $V_{L}$ is no longer a simple vertex algebra when $\det L=0$ in $R$. In particular, there may be quotients with exceptionally large symmetry. One concrete example is the $A_{2}$ lattice over $\mathbb{F}_{3}$, whose vertex algebra has a simple quotient with automorphism group $G_{2}(3)$: see [Griess-Lam 2013], Proposition 6.1 and Theorem 6.3  [^1] We thank C. H. Lam for bringing this example to our attention..
Viewed from a suitable distance, our proof is more or less a translation of Dong-Nagatomo’s proof, with some necessary adjustments. One such adjustment lies in the construction of $G_{L}$. Dong and Nagatomo construct $N$ by simply exponentiating the weight 1 Lie algebra of $V_{L,\mathbb{C}}$, but reductive algebraic groups in general cannot be constructed by exponentiating a Lie algebra. Instead, we construct $G_{L}$ by describing its torus $T=\operatorname{\underline{Hom}}(L,\mathbb{G}_{m})$ separately from its one-dimensional root subgroups $U_{\alpha}$, then showing that these subgroups satisfy the appropriate relations.
The main sticking point for us was the conjugacy of tori: over an algebraically closed field, all Cartan subgroups of a smooth affine group scheme are conjugate to each other, but this doesn’t hold for general commutative rings. The standard example is $SL_{2}(\mathbb{R})$, in which the diagonal torus is not conjugate to $SO(2)$. Since the maximal tori we consider are split, there are fairly strong replacement results: Proposition 6.2.11 in [Conrad 2014] asserts that if $G$ is split reductive over a nonempty connected scheme $S$ with trivial Picard group, then all split maximal tori are conjugate. It may be possible to turn this result into a solution, but we found it more straightforward to use étale-local conjugacy of all maximal tori. Using the étale sheaf property of affine schemes, we find that this is sufficient to show that all automorphisms are covered.
@@
In section 2, we introduce some basic theory of reductive group schemes, and define the group $G_{L}$. It is a split reductive group over $\mathbb{Z}$, and it acts faithfully on $V_{L}$.
In section 3, we define the finite flat group scheme O(L~)O($\tilde${L}), and show that it acts faithfully on $V_{L}$.
@@ Acknowledgements
S. C. would like to thank Robert Griess and Ching-Hung Lam for helpful discussions, and Arturo Pianzola for helpful comments on an earlier version of this paper. This work was supported by JSPS Kakenhi Grant-in-Aid for Young Scientists (B) 17K14152, and JSPS Kakenhi Grant-in-Aid for Scientific Research (C) 22K03264
@@ 1.1. Vertex algebras
    $u_{m},v_{n}]=\sum_{k\geq 0}\binom{m}{k}(u_{k}v)_{m+n-k},$
    or its equivalent $u_{m},Y(v,z)]=\sum_{k\geq 0}\binom{m}{k}Y(u_{k}v,z)z^{m-k}$
@@
  • (translation-covariance) $T^{(m)},Y(u,z)]=\sum_{i=1}^{m}\partial^{(i)}Y(u,z)T^{(m-i)}$, where $T^{(m)}:V\to V$ is the linear map $u\mapsto u_{-n-1}\mathbf{1}$.
@@
    $L_{0}$ acts diagonalizably with integer eigenvalues.
@@
    Each $V_{n}$ is a finite type projective $R$-module.
@@
    For all $m,n\in\mathbb{Z}$, $L_{m},L_{n}]=(m-n)L_{m+n}+\binom{m+1}{3}c\operatorname{id}_{V}$.
@@
The assumption that graded pieces are finite type and projective implies any finite generating set lies in some finite type projective module $U$ that is invariant under automorphisms, so we obtain an embedding in a product of general linear groups. It remains to show that the condition “vertex algebra multiplication is respected” is Zariski closed, so $\operatorname{\underline{Aut}}_{V/R}$ is represented by a closed affine group subscheme of some general linear group.
To prove the Zariski closed condition, we adapt the following variant of the argument in section 2 of [Dong-Griess 2001]. Let $U$ be a finite dimensional sum of homogeneous pieces of $V$ that contains a full generating set. For each $n$-tuple $(a_{1},\ldots,a_{n})$ of integers, we obtain a linear map $U^{\otimes n+1}\to V$ by composing the corresponding vertex operators:
$u^{1}\otimes\cdots\otimes u_{n+1}\mapsto u^{1}_{a_{1}}u^{2}_{a_{2}}\cdots u^{n}_{a_{n}}u^{n+1}$
For a homogeneous linear transformation $g$ on $U$ to be the restriction of an automorphism of $V$, it is necessary and sufficient that all of these products be preserved, for all $n\in\mathbb{Z}_{\geq 0}$. That is, for any $v\in V$, and any two presentations of $v$ as a finite sum of elements of the form $u^{1}_{a_{1}}u^{2}_{a_{2}}\cdots u^{n}_{a_{n}}u^{n+1}$, we require that the corresponding finite sums of elements of the form $g(u^{1})_{a_{1}}g(u^{2})_{a_{2}}\cdots g(u^{n})_{a_{n}}g(u^{n+1})$ coincide. Each pair of presentations of a vector $v$ thus yields a polynomial relation on the coordinates of $g$ (viewed as a linear transformation on $U$), and if $g$ satisfies all such relations, then it is the restriction of a uniquely defined automorphism of $V$. These polynomial relations therefore describe Zariski-closed subsets of $GL(U)$. ∎
@@ 1.2. Lattice vertex algebras
Let $L$ be an integer lattice, i.e., a finitely generated free abelian group equipped with a nondegenerate integer-valued symmetric bilinear form. We define the rational Heisenberg Lie algebra $H_{L}$ on $L$ to be the $\mathbb{Q}$-vector space $(L\otimes\mathbb{Q})[t,t^{-1}]\oplus\mathbb{Q}K$ with bracket given by $t^{i},t^{j}]=i\delta_{i+j,0}$ and $t_{i},K]=0$. This is a central extension of the abelian Lie algebra $(L\otimes\mathbb{Q})[t,t^{-1}]$ by the one dimensional Lie algebra $\mathbb{Q}K$. We write $H_{L}^{\geq 0}$ for the abelian subalgebra $(L\otimes\mathbb{Q})[t]\oplus\mathbb{C}K$, and for any $a\in L\otimes\mathbb{Q}$ and $r\in\mathbb{Q}$, we write $\mathbb{Q}_{a,r}$ for the 1-dimensional representation on which the subalgebra $t(L\otimes\mathbb{Q})[t]$ acts trivially, $(L\otimes\mathbb{Q})$ acts by inner product with $a$, and $K$ acts by the scalar $r$. Then, we write $\pi^{L}_{a,\mathbb{Q}}$ for the induced module $\operatorname{Ind}_{H_{L}^{\geq 0}}^{H_{L}}\mathbb{Q}_{a,1}$, and identify this space with $\operatorname{Sym}_{\mathbb{Q}}(t^{-1}(L\otimes\mathbb{Q})[t^{-1}])\otimes\mathbb{Q}e_{a}$ for a distinguished vector $e_{a}$. For any $b\in L\otimes\mathbb{Q}$ and $n\in\mathbb{Z}$, we define $b(n)$ to be the operator
@@
([Frenkel-Lepowsky-Meurman 1988] chapter 5) Let $L$ be an even lattice, i.e., an integer lattice where $(a,a)\in 2\mathbb{Z}$ for all $a\in L$. We define the double cover $\hat{L}$ to be the unique (up to isomorphism) central extension of $L$ by $\pm 1$ such that the commutator $[a,b]$ is given by $(-1)^{(\bar{a},\bar{b})}$, where $\bar{a},\bar{b}\in L$ are the images of $a,b\in\hat{L}$ under the covering homomorphism $\hat{L}\to L$. We will write $e:L\to\hat{L}$, given by $\alpha\mapsto e_{\alpha}$, to denote a set-theoretic section of the covering homomorphism. The twisted group ring $\mathbb{Z}\{L\}$ of $L$ is the quotient of $\mathbb{Z}[\hat{L}$] by the ideal generated by $\kappa+1$, where $\kappa$ is the nontrivial element in $\hat{L}$ that maps to $0\in L$. We write $^{L}\iota$$\hat$$\to\mathbb{Z}\{L\}$, given by $a\mapsto a\otimes 1$ for the multiplicative map. We note that there is an isomorphism $\mathbb{Z}[L]\to\mathbb{Z}\{L\}$ of abelian groups taking the basis vector $e^{\alpha}$ to $\iota(e_{\alpha})$, but this is in general not a ring homomorphism.
@@ 2.1. Fundamentals
We review the parts of the theory of reductive groups that are necessary for us. All definitions are taken from [SGA3]. A basic modern reference is section 1 of [Conrad 2014].
@@
By [SGA3] Exp. XIV Corollary 3.20, maximal tori of smooth group $R$-schemes exist Zariski locally on $\operatorname{Spec}R$. By [SGA3] Exp. XXII Corollary 5.3.11, any two maximal tori (of an fpqc locally split reductive group) are conjugate étale locally. This fact will play an essential role in our analysis of $\operatorname{Aut}V_{L}$.
@@
([SGA3] Exp. XIX Def. 1.6.1, 2.7) An affine group scheme over a field is reductive if it is smooth, geometrically connected, and its maximal connected unipotent normal subgroup is trivial (here, a group is unipotent if it admits an increasing filtration whose quotients are subgroups of the additive group $\mathbb{G}_{a}$, the group scheme which takes a ring to its underlying additive group). Given a commutative ring $R$, a reductive group scheme over $R$ is a group scheme $G\to\operatorname{Spec}R$ whose structure morphism is smooth and affine, and whose geometric fibers are reductive in the previous sense.
Remark 2.1.4.
There are other definitions of “reductive” in the literature, where fibers are not necessarily connected (e.g., [Conrad 2014]). We follow the definition in [SGA3] since that is where we draw the results we need.
@@
A root datum is reduced if for any root $a\in\Phi$, the only scalar multiples that are roots are a,$a$,-a. A root datum is semisimple if the saturation of $\mathbb{Z}\Phi$ in $X$ is all of $X$, or equivalently if $\mathbb{Q}\Phi=X\otimes\mathbb{Q}$.
@@
By [SGA3] Exp. XIV Corollary 3.20, for any reductive group scheme $G$ over a commutative ring $R$, and any point $x\in\operatorname{Spec}R$, there is a Zariski open neighborhood $U$ of $x$ such that $G_{U}$ has a maximal $U$-torus. Since such a torus is étale-locally split, we consider triples $(G,T,M)$, where $T=D_{R}(M)$ is split. Then, the adjoint $T$-action on the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ endows the vector bundle $\mathfrak{g}$ with an $M$-grading, splitting into a direct sum of $\operatorname{Lie}(T)$ and line bundles $\mathfrak{g}_{a}$. The corresponding sections $a\in M$ are called roots of the Lie algebra.
@@
In particular, $G$ is smooth and affine over $\mathbb{Z}$ with connected fibers, and its Lie algebra has distinguished Cartan subalgebra $X$ and roots given by $\Phi$.
@@
(basic properties of our group) $G_{L}$ is a smooth affine group scheme over $\mathbb{Z}$ with connected fibers. The Lie algebra of $G_{L}$ is the Lie algebra over $\mathbb{Z}$ given by the root datum $(X,\Phi,X^{\vee},\Phi^{\vee})$, and in particular is isomorphic to the weight 1 subspace $(V_{L})_{1}$, where the Cartan subalgebra $X$ is identified with the weight 1 subspace of $\pi^{L}_{0}$, with roots given by $\Phi$. The semisimple group $D(G_{L})$ is of ADE type (i.e., it is simply laced), and there is a central isogeny $Z\times D(G_{L})\to G_{L}$ with finite central kernel.
@@ 3. Symmetries of the Mewtwo cover
The symmetry group of this cover is then an extension O(L~)O($\tilde${L}) of $O(L)$ by $\operatorname{Hom}(L,\mu_{2})\cong\mu_{2}^{\operatorname{rank}L}$. We show that O(L~)O($\tilde${L}) acts faithfully on $V_{L}$.
@@ 3.2. Orthogonal transformations
Let $(\tilde{A},\pi)$ be the $\mu_{2}$-cover of a free abelian group $A$ with quadratic form $q:A\to\mathbb{Z}$. An orthogonal transformation of $(\tilde{A},\pi)$ over $S$ is a group scheme automorphism $\phi$ of $\tilde{A}_{S}$ such that there exists an automorphism $\bar{\phi}$ of $A$ that preserves $q$, satisfying $\pi\circ\phi=\bar{\phi}\circ\pi$. The functor that takes an $R$-algebra $S$ to the group of orthogonal transformations of $(\tilde{A},\pi)$ over $S$ is written O(A~)O($\tilde${A}), or $O(\tilde{A},\pi)$.
@@
[…]
@@
If $L$ is positive-definite, then O(L~)O($\tilde${L}) is an affine group scheme of finite type.
@@
Given a basis of $L$, let $d$ be the maximal norm of vectors in this basis, and let $n$ be the number of vectors in $L$ of length at most $d$. Then O(L~)O($\tilde${L}) is a subgroup of $GL_{n}$, since it acts faithfully on the linear space whose basis is the set of these vectors. ∎